# The Catalogue of Virtual Early-Type Galaxies from IllustrisTNG: Validation and Real Observation Consistency

### v1

@@ 3.4.1 Effective radius versus stellar half-mass radius
We start by studying the relation between $r_{1/2}$ and $R_{\rm e}$, summarized in Fig. 12. While $R_{\rm e}$ is the (projected) radius enclosing half of the total luminosity of a galaxy, $r_{1/2}$ is the radius of the sphere enclosing half of its total stellar mass, meaning that they are not the straight projected definition of the other. To be fully consistent, one should compare $R_{\rm e}$ to the 3D half-light radius. However, following the arguments in Wu et al. (2024), as the stellar mass-to-light ratio gradients in 3D are closer to unit than their 2D counterpart, we take the 3D stellar half-mass radius as being a reasonable approximation of the 3D half-light radius. We observe that galaxies associated with higher Sérsic index are typically the ones where $R_{\rm e}>r_{1/2}$. Besides these particular objects, the slope $\partial\log R_{\rm e}/\partial\log r_{1/2}=0.87$ and offset value of $-0.049$ shows that $R_{\rm e}$ is systematically smaller than $r_{1/2}$ for the virtual-ETG sample, as one should expect. The linear relationship between these two quantities is, therefore, expressed as:
@@
Using the fit provided in Lima Neto et al. (1999) for the relation between $r_{1/2}/R_{\rm e}$ and the reciprocal of the Sérsic index $n$, Wolf et al. (2010) have verified that $r_{1/2}/R_{\rm e}\approx 4/3$ is an accurate approximation for most surface brightness profiles used to model galaxies’ radial light distributions. By directly performing the ratio between $r_{1/2}$ and $R_{\rm e}$ for the galaxies in our catalogue, we have verified that the mean ratio is $\langle r_{1/2}/R_{\rm e}\rangle=1.39$ with a scatter $\sigma=0.23$, consistent with Wolf et al. (2010) statement. Typical values of $r_{1/2}/R_{\rm e}$ are also computed in Ciotti (1991) (see their table 2) and they vary from $1.34$ to $1.35$. For ATLAS3D ETGs, Cappellari et al. (2013a) found a mean ratio $\langle r_{1/2}/R_{\rm e}\rangle\approx 1.42$ which is $1\sigma$ consistent with the ratio we found. Overall, these results show agreement between real world samples of galaxies and our virtual-ETG sample for the $r_{1/2}-R_{\rm e}$ relation.
@@ 3.4.2 Projected versus three-dimensional luminosity
In practice, divergences can occur because theoretical predictions are only approximations to the fiducial results. Regarding the Fig. 13, we believe that differences between the expected behaviour and the theoretical expectations occur because of the properties of objects at the low- and the high-luminosity ends. At the high luminosity end, there is a greater chance that large galaxies have extra luminous features on their projected light distributions (e.g., diffuse light that were not completely ruled out by the $30$ kpc cut) and hence, are not generally described by the one-component Sérsic model. Despite that, as can be observed in Fig. 13, these objects represent only a tiny fraction and almost all galaxies have $L_{2D}>L_{3D}$ at this bright end. At the low luminosity end, there is a greater chance that smaller galaxies have more irregular (noisy) 2D profiles, leading to possible mismatches between the real 2D light and the Sérsic light. This is also the source of the larger scatter observed at the low luminosity end. To verify this assertion, we have plotted the median behaviour of the $\chi^{2}$ values for the Sérsic fits versus the 3D luminosity at both $1R_{\rm e}$ and $2R_{\rm e}$ apertures, presented in Fig. 14. It is clear from this Figure that $\chi^{2}$ drops as luminosity increases, indicating that fainter (smaller) galaxies have profiles that are not well described by the one-component Sérsic model used here.
@@ Appendix B Parameters and statistical properties of the catalogue
In this appendix, we present a table containing all the parameters in the catalogue (see Table 3). Some parameters as the total simulated stellar mass, specific star formation rate, etc., are not included because they can be easily derived from the quantities already included.
@@ Appendix C TNG100-1 velocity dispersion profiles
In Fig. 17, we show $142$ velocity dispersion profiles for the ETGs in our sample. For the innermost regions, i.e., below the softening length ($\epsilon_{\star}=0.74$ kpc; see the dashed black vertical line in the Figure), a nearly constant behaviour or a drop on the profiles instead of a central peak is observed, and galaxies with lower stellar mass are affected the most. As commented on the main text, this effect, is artificially induced by the smoothing of the gravitational field for scales below the simulation’s softening length (“softened velocity dispersion” for short) and contrasts the usually peaked velocity dispersion profiles of observed massive ETGs (see, e.g., Coccato et al. 2009; Cappellari et al. 2013a).

### v2

@@ 3.4.1 Effective radius versus stellar half-mass radius
We start by studying the relation between $r_{1/2}$ and $R_{\rm e}$, summarized in Fig. 11. While $R_{\rm e}$ is the (projected) radius enclosing half of the total luminosity of a galaxy, $r_{1/2}$ is the radius of the sphere enclosing half of its total stellar mass, meaning that is not the straight 3D definition of $R_{\rm e}$. To be fully consistent, one should compare $R_{\rm e}$ to the 3D half-light radius. However, following the arguments in Wu et al. (2024), as the stellar mass-to-light ratio gradients in 3D are closer to unit than their 2D counterpart, we take the 3D stellar half-mass radius as being a reasonable approximation of the 3D half-light radius. We observe that galaxies associated with higher Sérsic index are typically the ones where $R_{\rm e}>r_{1/2}$. Besides these particular objects, the slope $\partial\log R_{\rm e}/\partial\log r_{1/2}=0.87$ and offset value of $-0.049$ shows that $R_{\rm e}$ is systematically smaller than $r_{1/2}$ for the virtual-ETG sample, as one should expect. The linear relationship between these two quantities is, therefore, expressed as:
@@
Using the fit provided in Lima Neto et al. (1999) for the relation between $r_{1/2}/R_{\rm e}$ and the reciprocal of the Sérsic index $n$, Wolf et al. (2010) verified that for most surface brightness profiles used to model galaxy light distributions, the ratio $r_{1/2}/R_{\rm e}\approx 4/3$ is an accurate approximation regardless the profiles’ shapes. By directly performing the ratio between $r_{1/2}$ and $R_{\rm e}$ for the galaxies in our catalogue, we have verified that the mean ratio is $\langle r_{1/2}/R_{\rm e}\rangle=1.39$ with a scatter $\sigma=0.23$, consistent with Wolf et al. (2010) statement. Typical values of $r_{1/2}/R_{\rm e}$ are also computed in Ciotti (1991) (see their table 2) and they vary from $1.34$ to $1.35$. For ATLAS3D ETGs, Cappellari et al. (2013a) found a mean ratio $\langle r_{1/2}/R_{\rm e}\rangle\approx 1.42$ which is $1\sigma$ consistent with the ratio we found. Overall, these results show agreement between real world samples of galaxies and our virtual-ETG sample for the $r_{1/2}-R_{\rm e}$ relation.
@@ 3.4.2 Projected versus three-dimensional luminosity
Regarding the Fig. 12, we believe that differences between the expected behaviour and the theoretical expectations occur because of the properties of objects at the low- and the high-luminosity ends. At the high luminosity end, there is a greater chance that large galaxies have extra luminous features on their projected light distributions (e.g., diffuse light that were not completely ruled out by the $30$ kpc cut) and hence, are not generally described by the one-component Sérsic model. Despite that, as can be observed in Fig. 12, these objects represent only a tiny fraction and almost all galaxies have $L_{2D}>L_{3D}$ at this bright end. At the low luminosity end, there is a greater chance that smaller galaxies have more irregular (noisy) 2D profiles, leading to possible mismatches between the real 2D light and the Sérsic light. This is also the source of the larger scatter observed at the low luminosity end. To verify this assertion, we have plotted the median behaviour of the $\chi^{2}$ values for the Sérsic fits versus the 3D luminosity at both $1R_{\rm e}$ and $2R_{\rm e}$ apertures, presented in Fig. 13. It is clear from this Figure that $\chi^{2}$ drops as luminosity increases, indicating that fainter (smaller) galaxies have profiles that are not well described by the one-component Sérsic model used here.
@@ Appendix B Parameters and statistical properties of the catalogue
In this appendix, we present a table containing all the parameters in the catalogue (see Table 3). Some parameters as the specific star formation rate, are not included because they can be easily derived from the quantities already included.
@@ Appendix C TNG100-1 velocity dispersion profiles
In Fig. 16, we show $142$ velocity dispersion profiles for the ETGs in our sample. For the innermost regions, i.e., below the softening length ($\epsilon_{\star}=0.74$ kpc; see the dashed black vertical line in the Figure), a nearly constant behaviour or a drop on the profiles instead of a central peak is observed, and galaxies with lower stellar mass are affected the most. As commented on the main text, this effect, is artificially induced by the smoothing of the gravitational field for scales below the simulation’s softening length (“softened velocity dispersion” for short) and contrasts the usually peaked velocity dieprsion profiles of observed massive ETGs (see, e.g., Coccato et al. 2009; Cappellari et al. 2013a).